# WallFacer: Harnessing Multi-dimensional Ring Parallelism for Efficient Long Sequence Model Training

### v3

@@ Abstract
Training Transformer models on long sequences in a distributed setting poses significant challenges in terms of efficiency and scalability. Current methods are either constrained by the number of attention heads or excessive communication overheads. To address this problem, we propose WallFacer, a multi-dimensional distributed training system for long sequences, fostering an efficient communication paradigm and providing additional tuning flexibility for communication arrangements. Specifically, WallFacer introduces an extra parallel dimension to substantially reduce communication volume and avoid bandwidth bottlenecks. Through comprehensive experiments across diverse hardware environments and on both Natural Language Processing (NLP) and Computer Vision (CV) tasks, we demonstrate that our approach significantly surpasses state-of-the-art methods that support near-infinite sequence lengths, achieving performance improvements of up to 77.12% on GPT-style models and up to 114.33% on DiT (Diffusion Transformer) models.
@@ 1 Introduction
Over the past decade, Transformer[vaswani2023attention] models have made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP). As the technology has evolved, the ability to efficiently process long sequences with Transformer has emerged as a pivotal challenge. For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books [Koh2022AnES, beltagy2020longformer]. Similarly, chat-based applications, such as ChatGPT [openai2023gpt4], require the capacity to process extensive dialogue histories to ensure conversational consistency. There are also applications in other fields like video generation[Brooks_2024, peebles2023scalable] and protein structure prediction[Jumper2021HighlyAP, cheng2023fastfold].
Figure 1: WallFacer-2 and WallFacer-4 theoretically save around 50% and 75% of total P2P communication volume for various sequence length.
@@
Traditional parallelisms such as Data Parallelism[hillis1986data, shoeybi2019megatron, li2020taming, you2018imagenet], Tensor Parallelism[shoeybi2019megatron, wang2022tesseract, xu2021efficient], and Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] distribute the model, input batch, and the optimizer states, but can not directly address the large memory requirement of extremely long sequences as the sequence length dimension remains unchanged. To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension. Mainstream Sequence Parallelism schemes can generally be classified into two categories: those based on all-to-all communication, and those based on ring peer-to-peer communication. Methods like DeepSpeed Ulysses[jacobs2023Ulysses], which are based on all-to-all communication, offer efficiency but require the splitting of attention heads. Consequently, these methods are limited in scalability and can not be scaled to more devices than the number of attention heads. On the other hand, peer-to-peer communication methods[liu2023ring, li2024distflashattn], such as Ring Attention[liu2023ring], do allow for near-infinite context lengths; however, they require the transmission of complete keys and values across all GPUs, leading to significantly high communication loads. In summary, there remains a deficiency in communication-efficient methods that are capable of supporting near-infinite context lengths.
Table 1: Meanings of the symbols that are used in this paper
@@
$C$ The parallel size of WallFacer (attention parallel size)
@@
$N$ The total number of tokens within the whole sequence
@@
Inspired by the methodology of n-body communication optimization [Driscoll2013ACN], we introduce WallFacer, a novel near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension into the existing ring-style communication. Specifically, instead of including all GPUs in a single parallel group as done in Ring Attention [liu2023ring], WallFacer groups the GPUs into teams and divides the peer-to-peer communication within these teams. This approach fosters an efficient communication paradigm and provides extra tuning flexibility for communication arrangements. With very little additional memory cost, WallFacer parallelism significantly reduces the peer-to-peer communication volume, as shown in Figure 1. Compared to previous works, WallFacer is not limited in supported sequence length by attention heads like DeepSpeed Ulysses[jacobs2023Ulysses] and Megatron Sequence Parallelism[korthikanti2022reducing], and also shows better communication efficiency and scalability than Ring Attention[liu2023ring].
In summary, our paper presents these contributions:
  • We introduce a near-infinite-context training system for Transformer models, featuring a groundbreaking multi-ring sequence parallelism scheme. This scheme adds an additional dimension of parallelism and significantly reduces peer-to-peer communication, all while maintaining a minimal memory footprint.
  • We offer a straightforward grid-search method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
  • We perform experiments on mainstream Transformer models, including GPT-style[gpt] and DiT-style[peebles2023scalablediffusionmodelstransformers], conducting performance and scaling tests across various computing clusters. Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12% on the GPT model and up to 114.33% on the DiT model, showcasing its efficacy and scalability.
2 Background and Motivation
2.1 Transformer models and Long Sequence Training
Transformer-based models have become the mainstream approach in both Natural Language Processing (NLP) and Computer Vision (CV). GPT-style models [gpt] lead the field in NLP tasks, generating tokens in an auto-regressive manner. Meanwhile, the Diffusion Transformer (DiT) [peebles2023scalablediffusionmodelstransformers], which employs pure Transformer blocks within traditional diffusion processes, has gained widespread use over the past year for image and video generation. The key mechanism behind these Transformer-based models is attention[vaswani2023attention], which captures the text feature by calculating the attention score between every two single tokens. A standard attention function is given as:
$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
where $Q$, $K$, and $V$ are the query, value, and key of input x with shape ($B$, $N$, $H$) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention. For convenience, we use the symbolic representations in Table 1. We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length. To solve the memory IO and capacity bottleneck, methods like flash-attention[dao2022flashattention] have been proposed. Flash-attention is designed upon the idea of softmax decomposition [kitaev2020reformer, milakov2018online, rabe2022selfattention], enabling online softmax computation by storing some extra statistics. Flash-attention is thus capable of segmenting the $Q$, $K$, and $V$ matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially. This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM).
However, the sequence length can reach hundreds of thousands, when dealing with multi-round chatting, or high-resolution long video generation. It then becomes necessary to distribute the sequence across multiple GPUs. This distribution helps to reduce both the memory and computation demands on any single device. This strategy is also known as sequence parallelism. We will explore these methods in further detail in the following section.
2.2 Sequence Parallelism
Sequence parallelism was first conceptualized in a study published in 2021 [Li2021SequencePL], which introduced the concept as a strategy to distribute computational sequences across multiple processing steps. Presently, Sequence parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based. The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of flash-attention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics.
2.2.1 Attention-Head-Sharding-Based
Here we introduce the two representative methods, DeepSpeed Ulysses and Megatron Sequence Parallelism.
DeepSpeed Ulysses. DeepSpeed Ulysses[jacobs2023Ulysses] presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability. Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism with two all-to-all communication. It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure.
The principal merit of Ulysses lies in its simplicity and ease of implementation. Nonetheless, its reliance on the number of attention heads for partitioning activations introduces the following limitations. 1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which restricts the maximum sequence length that can be processed. This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) [ainslie2023gqa] or multi-query attention (MQA) [shazeer2019fast], which further reduce the number of available attention heads for Keys and Values. 2) Ulysses encounters load-balancing issues when the number of attention heads is not evenly divisible by the number of GPUs, complicating its deployment across diverse hardware configurations. 3) The need for head-splitting makes Ulysses hard to combine with Tensor Parallelism.
Megatron Sequence Parallelism. Megatron Sequence Parallelism focuses on minimizing memory usage and reducing the necessity for activation recomputation. The training approach is based on Tensor Parallelism (TP). While TP is employed in the Linear and Self-Attention blocks, operations such as the LayerNorm and Dropout functions are not distributed, leading to replications within the tensor parallel group. This requires an additional all-gather operation and a further reduce-scatter operation during the forward propagation to alternate between Tensor Parallelism and Sequence Parallelism. The Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length. Like DeepSpeed Ulysses, it is also limited by the number of attention heads.
2.2.2 Ring-peer-to-peer-communication-based
The primary method in peer-to-peer-communication-based strategies is Ring Attention[liu2023ring].
Ring Attention. Introduced in 2023, Ring Attention[liu2023ring] innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs. Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2. This method employs an online-softmax and updates attention scores incrementally, allowing the computation of attention scores without retaining the full sequence length. Thus, it potentially supports infinite context, provided sufficient computing resources are available. However, the requirement for $P$ steps of P2P communication renders this approach less efficient in environments with high-latency communication.
Figure 3: An overview of the WallFacer Training System
2.3 Motivation
Through observation, we identify two main drawbacks of Ring Attention. First, the communication overhead is exceedingly high because every GPU in the system must send and receive keys and values for nearly the entire sequence length before completing the attention computation. Second, variations in bandwidth between and within computing nodes can cause communication bottlenecks. As illustrated in Figure 2, the bandwidths between GPUs 3 and 4 and between GPUs 11 and 12 are lower than those between other GPUs in the ring. Despite this, the system is forced to operate in a complete circle, which can result in unnecessary idle times for other GPUs.
Inspired by techniques used in n-body simulations [Hockney_Eastwood_1989], we propose an approach that addresses both challenges simultaneously. In n-body simulations, such as those calculating direct forces between particles (e.g., gravitational forces), researchers discovered that communication overhead can be minimized by introducing additional dimensions of communication [Driscoll2013ACN]. Similarly, we propose grouping GPUs into teams within the ring-style sequence parallelism framework and distributing communication tasks among team members. This approach reduces the overall communication volume and mitigates the bandwidth bottlenecks previously mentioned. Building upon this concept, we developed the WallFacer training system, which we will detail in the following section.
3 WallFacer Training System
3.1 Overview
The WallFacer training system is comprised of five main components. At the core of the system is WallFacer Attention, which leverages multiple ring-style P2P (peer-to-peer) communication strategies to enhance the efficiency of distributed attention computation. The Communication Configuration Generator plays a critical role in initially assigning keys and values to their corresponding ranks. Meanwhile, the Communication Topology Scheduler outlines the placement of parallelism across various computing nodes. The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs. Finally, the WallFacer Runtime incorporates additional supporting techniques for the training process, such as gradient checkpointing. We will now explore the details of these components in depth.
3.2 WallFacer Attention
Figure 4: An example of WallFacer Attention on 64 GPUs. Each team member forms a sub-ring with members with the same local rank from other teams in the same ring communication group, reducing the communication volume of each team member by 75%.
As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes. We enhance the ring sequence parallelism by introducing an additional dimension. This is achieved by duplicating the queries, keys, and values, thus dividing the communication tasks and distributing them within each team. This segmentation markedly improves computational efficiency and scalability.
We will now go through the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
3.2.1 Training Process of WallFacer Parallelism
@@
In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more efficiently, unlike the isolated operations in traditional settings. The queries, keys, and values are gathered within the team, so each member of the team holds the same activation of the whole team before attention. WallFacer introduces an additional parameter, $C$, which determines the replication factor of the input and, consequently, the number of GPUs within each team. The range of $C$ is from 1 to $\sqrt{P}$. When $C$ equals one, the algorithm falls back to Ring Attention. When $C$ equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one with no rings. When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently. There are three main distinctions between the operational processes of WallFacer and Ring Attention: QKV_matmul & all-gather, attention iteration, and reduce-scatter. Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks. The complete forward process of a WallFacer Transformer block is outlined in Algorithm 1.
Algorithm 1 WallFacer Attention Block (Forward)
@@
4:  initialize attention score O, extra statistics lse to zero.
5:  for $1\leq i\leq$world_size$/c^{2}$ do
@@
Forward Propagation. In Figure 4, we have an example of one team of four GPUs out of all the 64 GPUs performing WallFacer-style attention. Each training iteration begins with the dataloader splitting the entire input sequence of length $N$ into $N/P$ sub-sequences, which are then loaded onto each GPU. As previously mentioned, the next step involves computing the queries, keys, and values. These are computed separately via matrix multiplication, followed immediately by the launch of the all-gather kernel, which gathers the above QKVs within the team, allowing for the overlap of up to two-thirds of the communication with computation.
@@
Once this phase is complete, each GPU within the team possesses the same Q, K, and V, each of a length of $\frac{CN}{P}$. To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU by the Communication Configuration Generator. These ranks determine each GPU’s partners and position within the P2P ring, details of which will be discussed in Section 3.3.
@@
Following the setup, the Keys and Values are dispatched to their designated locations within the cluster to establish the initial sub-ring, setting the stage for the multi-ring iteration phase of WallFacer attention. Given that each sub-sequence is $\frac{CN}{P}$ long and each GPU is tasked with computing the attention score for $\frac{1}{C}$ of the whole sequence, it results in $\frac{N/C}{CN/P}=P/C^{2}$ rounds of communication. This implies that there are $P/C^{2}$ GPUs in one ring.
@@
The iteration process, conducted in a Flash-Attention style, involves storing the log-sum-exp (lse) and intermediate output O, which are updated step by step. Queries are retained locally, while Keys and Values circulate through the ring via P2P communication. After completing the iterations, each team member accumulates the attention scores for the entire team’s sub-sequence of Queries with $1/C$ of the Keys and Values from the full sequence.
@@
A simple reduce-scatter operation is then employed to amalgamate the intermediate results and distribute them among the team members. Each GPU ultimately contains the final attention score for its portion of the sequence over the entire sequence. The output of this forward attention block is finalized after a standard LayerNorm and FeedForward layer process.
Backward Propagation. The major distinction between backward and forward propagation is the inability to calculate queries independently during the backward phase. Unlike forward propagation, the backward phase requires the complete set of keys and values to calculate the gradient for queries, and vice versa. To manage this, we have structured the gradient calculation into two loops: the key & value outer loop and the query inner loop.
@@
In the outer loop, gradients for keys and values are tracked and maintained fixed on the corresponding GPUs within the sub-rings; these gradients do not transfer between GPUs. The inner loop, however, handles the gradients for queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves. During each iteration, the approach mirrors the backward computation method used in flash-attention[dao2022flashattention], where the updated gradient of the current query shard is passed to the next GPU in the ring, while the gradients for keys and Values are retained for subsequent query shards.
Initial communication ranks for sending and receiving are still configured by the Communication Configuration Generator. Additionally, there is an extra P2P communication required at the end of the process to send the query and its gradient back to their original location following the loop.

### v4

# StarTrail: Concentric Ring Sequence Parallelism for Efficient Near-Infinite-Context Transformer Model Training

@@ Abstract
Training Transformer models on long sequences in a distributed setting poses significant challenges in terms of efficiency and scalability. Current methods are either constrained by the number of attention heads or excessive communication overheads. To address this problem, we propose StarTrail, a multi-dimensional concentric distributed training system for long sequences, fostering an efficient communication paradigm and providing additional tuning flexibility for communication arrangements. Specifically, StarTrail introduces an extra parallel dimension and divides the peer-to-peer communication into sub-rings to substantially reduce communication volume and avoid bandwidth bottlenecks. Through comprehensive experiments across diverse hardware environments and on both Natural Language Processing (NLP) and Computer Vision (CV) tasks, we demonstrate that our approach significantly surpasses state-of-the-art methods that support Long sequence lengths, achieving performance improvements of up to 77.12% on GPT-style models and up to 114.33% on DiT (Diffusion Transformer) models without affecting the computation results.
@@ 1 Introduction
Over the past decade, Transformervaswani2023attention models have made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP). As the technology has evolved, the ability to efficiently process long sequences with Transformer has emerged as a pivotal challenge. For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books Koh2022AnES ; beltagy2020longformer . Similarly, chat-based applications, such as ChatGPT openai2023gpt4 , require the capacity to process extensive dialogue histories to ensure conversational consistency. There are also applications in other fields like video generationBrooks_2024 ; peebles2023scalable and protein structure predictionJumper2021HighlyAP ; cheng2023fastfold .
@@
Figure 1: StarTrail-2 and StarTrail-4 theoretically save around 50% and 75% of total P2P communication volume for various sequence length and achieves up to 2x speedup in end-to-end training
@@
Traditional parallelisms such as Data Parallelismhillis1986data ; shoeybi2019megatron ; li2020taming ; you2018imagenet , Tensor Parallelismshoeybi2019megatron ; wang2022tesseract ; xu2021efficient , and Pipeline Parallelismgpipe ; dapple ; li2021chimera ; Liu2023Hanayo distribute the model, input batch, and the optimizer states, but can not directly address the large memory requirement of extremely long sequences as the sequence length dimension remains unchanged. To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension. Mainstream Sequence Parallelism schemes can generally be classified into two categories, and are usually combined fang2024uspunifiedsequenceparallelism to complement each other’s drawbacks. Methods like DeepSpeed Ulyssesjacobs2023Ulysses , which are based on all-to-all communication, offer efficiency but require the splitting of attention heads. Consequently, these methods are limited in scalability and can not be scaled to more devices than the number of attention heads. On the other hand, peer-to-peer communication methodsliu2023ring ; li2024distflashattn , such as Ring Attentionliu2023ring , do allow for near-infinite context lengths; however, they require the transmission of complete keys and values across all GPUs, leading to significantly high communication loads. In summary, there remains a deficiency in communication-efficient methods that are capable of supporting near-infinite context lengths. In this paper, we focus on solving the communication inefficiency of ring-style sequence parallelism.
To solve these challenges, we introduce StarTrail, a novel near-infinite-context Transformer training system with concentric multi-ring sequence parallelism that incorporates an additional parallel dimension into the existing ring-style communication. Specifically, instead of including all GPUs in a single parallel group as done in Ring Attention liu2023ring , StarTrail groups the GPUs into teams and divides the peer-to-peer communication within these teams. This approach fosters an efficient communication paradigm and provides extra tuning flexibility for communication arrangements. With very little additional memory cost, StarTrail parallelism significantly reduces the peer-to-peer communication volume, as shown in Figure 1. Compared to previous works, StarTrail is not limited in supported sequence length by attention heads like DeepSpeed Ulyssesjacobs2023Ulysses and Megatron Sequence Parallelismkorthikanti2022reducing , and also shows better communication efficiency and scalability than Ring Attentionliu2023ring . We perform experiments on mainstream Transformer models, including GPT-stylegpt and DiT-stylepeebles2023scalablediffusionmodelstransformers , conducting performance and scaling tests across various computing clusters. Experiment results indicate that our StarTrail system outperforms Ring Attention by up to 77.12% on the GPT model and up to 114.33% on the DiT model, showcasing its efficiency and scalability.
2 Background and Related Works
2.1 Long Sequence Training and Sequence Parallelism
The key mechanism behind these Transformer-based models is attentionvaswani2023attention , which captures the text feature by calculating the attention score between every two single tokens. However, the sequence length can reach hundreds of thousands, when dealing with multi-round chatting, or high-resolution long video generation. It then becomes necessary to distribute the sequence across multiple GPUs. This distribution helps to reduce both the memory and computation demands on any single device. This strategy is also known as sequence parallelism. Presently, Sequence parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based. The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of FlashAttention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics.
2.1.1 Ring-peer-to-peer-communication-based
The primary method in peer-to-peer-communication-based strategies is Ring Attentionliu2023ring , which is also the main baseline of this work. Introduced in 2023, Ring Attentionliu2023ring innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs. Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2. This method employs an online-softmax and updates attention scores incrementally, allowing the computation of attention scores without retaining the full sequence length. Thus, it potentially supports infinite context, provided sufficient computing resources are available. However, the requirement for the same number of rounds of P2P communication as the number of GPUs renders this approach less efficient in environments with high-latency communication.
Figure 3: An overview of the StarTrail Training System
2.1.2 Attention-Head-Sharding-Based
There are two representative methods, DeepSpeed-Ulysses and Megatron Sequence Parallelism in this category. DeepSpeed-Ulyssesjacobs2023Ulysses transitions from sequence parallelism to a method akin to tensor parallelism with two all-to-all communication. It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure. Megatron Sequence Parallelismkorthikanti2022reducing focuses on minimizing memory usage and reducing the necessity for activation recomputation rather than efficiency. The two methods both rely on the number of attention heads to split the sequence, thus limited in scalability, especially when employing techniques like grouped-query attention (GQA) ainslie2023gqa or multi-query attention (MQA) shazeer2019fast . As these two sequence parallelism methods are orthogonal to the ring-based method, they are usually combined with ring attention to enable longer sequences.
In this paper, we focus on optimizing ring-style sequence parallelism, noting that integrating it with DeepSpeed-Ulysses Parallelism does not interfere with the underlying ring process.
3 StarTrail Training System
Figure 4: An example of StarTrail Attention on 64 GPUs. Each team member forms a sub-ring with members with the same local rank from other teams in the same ring communication group, reducing the communication volume of each team member by 75%.
3.1 Motivation
Through observation, we identify two main drawbacks of Ring Attention. First, the communication overhead is exceedingly high because every GPU in the system must send and receive keys and values for nearly the entire sequence length before completing the attention computation. Second, variations in bandwidth between and within computing nodes can cause communication bottlenecks. As illustrated in Figure 2, the bandwidths between GPUs 3 and 4 and between GPUs 11 and 12 are lower than those between other GPUs in the ring. Despite this, the system is forced to operate in a complete circle, which can result in unnecessary idle times for other GPUs. To solve these drawbacks, we develop the StarTrail training system, which we will detail in the following section.
3.2 StarTrail Attention
Figure 5: Meanings of the symbols that are used in this paper
@@ 3.2 StarTrail Attention
$C$ The parallel size of StarTrail (team size)
@@
NN The total number of tokens within the whole sequence
@@
As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes. We enhance the ring sequence parallelism by introducing an additional dimension. The fundamental idea of StarTrail is akin to a divide-and-conquer strategy. During attention, each token must compute its attention score with every other token in the sequence. While Ring Attention passes keys and values along a ring of $P$ GPUs over P1P-1 iterations, our approach introduces the concept of a team. In this setting, each team member interacts only with a designated portion of the overall sequence, and the results are later aggregated using collective communication. Thus, StarTrail can be devided into three phases: preprocessing, ring-phase, and postprocessing.
For preprocessing, we duplicate the queries within a team using an all-gather operation. This ensures that when a GPU receives new keys and values, it can compute the attention scores for the entire team’s queries. Similarly, gathering the keys and values allows us to reduce the number of P2P communication iterations by transmitting longer sequences per iteration. Following the preprocessing, we enter the ring-style communication phase. With the number of GPUs in one team being $C$, $CN/P$ tokens are exchanged in each iteration, and each GPU is responsible for computing $N/C$ tokens. This leads to a number of iterations of $\frac{N/C}{CN/P}=\frac{P}{C^{2}},$ within a smaller ring, which we refer to as a subring. For convenience, we group $\frac{P}{C^{2}}$ adjacent teams into a team group for subring communication, where GPUs sharing the same local team rank form the subring. An initial P2P communication step is executed to ensure that each team group has access to the complete set of keys and values for the sequence (details are provided in the Appendix). After completing the subring iterations, each GPU holds $1/C$ of the overall computation result for its team. With the help of online softmax, we then apply a simple reduce-scatter to combine these results while eliminating the duplicate tokens, which we refer to as the postprocessing. Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks. Now we will delve into more details in the StarTrail training process.
3.2.1 Configurations of StarTrail Parallelism
@@
In the StarTrail system, GPUs are grouped into Teams to coordinate computation and communication tasks more efficiently. StarTrail introduces an additional parameter, $C$, which determines the replication factor of the input and, consequently, the number of GPUs within each team. The range of $C$ is from 1 to $\sqrt{P}$. When $C$ equals one, the algorithm falls back to Ring Attention. When $C$ equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one with no rings. When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently.
Algorithm 1 StarTrail Attention Block (Forward)
@@
4: initialize attention score O, extra statistics lse to zero. // lse stands for log-sum-exp
[…]
@@
Forward Propagation. In Figure 4, we have an example of one team of four GPUs out of all the 64 GPUs performing StarTrail-style attention. Each training iteration begins with the dataloader splitting the entire input sequence of length $N$ into $N/P$ sub-sequences, which are then loaded onto each GPU. As previously mentioned, the next step involves computing the queries, keys, and values. These are computed separately via matrix multiplication, followed immediately by the launch of the all-gather kernel, which gathers the above QKVs within the team, allowing for the overlap of up to two-thirds of the communication with computation.
Figure 6: An example of ring initialization process of 8GPUs and 4 sub-rings in StarTrail.
@@
Once this phase is complete, each GPU within the team possesses the same Q, K, and V, each of a length of $\frac{CN}{P}$. To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU. These ranks determine each GPU’s partners and position within the P2P ring, as is shown in Figure 6.
@@
Following the setup, the Keys and Values are dispatched to their designated locations within the cluster to establish the initial sub-ring, setting the stage for the multi-ring iteration phase of StarTrail attention. Given that each sub-sequence is $\frac{CN}{P}$ long and each GPU is tasked with computing the attention score for $\frac{1}{C}$ of the whole sequence, it results in $\frac{N/C}{CN/P}-1=P/C^{2}-1$ rounds of communication. This implies that there are $P/C^{2}$ GPUs in one ring.
@@
The iteration process involves storing the log-sum-exp (lse) and intermediate output O, which are updated step by step. Queries are retained locally, while Keys and Values circulate through the ring via P2P communication. After completing the iterations, each team member accumulates the attention scores for the entire team’s sub-sequence of Queries with $1/C$ of the Keys and Values from the full sequence.
@@
A simple reduce-scatter operation is then employed to amalgamate the intermediate results and distribute them among the team members. Each GPU ultimately contains the final attention score for its portion of the sequence over the entire sequence.
@@
Backward Propagation. The major distinction between backward and forward propagation is the inability to calculate queries independently during the backward phase. Unlike forward propagation, the backward phase requires the complete set of keys and values to calculate the gradient for queries, and vice versa. To manage this, we have structured the gradient calculation into two loops: the key & value outer loop and the query inner loop. In the outer loop, gradients for keys and values are tracked and maintained fixed on the corresponding GPUs within the sub-rings; these gradients do not transfer between GPUs. The inner loop, however, handles the gradients for queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves. During each iteration, the approach mirrors the backward computation method used in FlashAttentiondao2022flashattention , where the updated gradient of the current query shard is passed to the next GPU in the ring, while the gradients for keys and Values are retained for subsequent query shards.